User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Simin

Attended: Jamie, Mas, Katie, Lucas, Tiege, Atifur, Jesse, Prad


Absent: Ross, Patrick, Arwin

Handover Issues

Katherine 12m:

  • Still have problem with Clock jump/delay drift.

For the moment try to shut the dbbc off and then bring it back. Jamie still working on it. One jump is ok in one hour but if it is more than one, it is better to restart it.

Hobart 12m:

  • Antenna was stuck and Lucas had to fix it(aum009).

Sometimes when the antenna is turned off and the field system is running, it loses the connection so sometimes it is needed to reboot the drive control. Before going to the antenna check some buttons and use wiki commands.

Yarragadee 12m:

  • Drive full issue. It seems monica files occupy a lot of space. Jamie will keep eye on it to make more space.

General issues:

  • Sound alarm: If you couldn't hear the sound, first check the black box behind the monitors to check if it is on or not.
/home/www/auscope/opswiki/data/pages/operations/agenda2018_09_20.txt · Last modified: 2018/09/20 04:23 by Simin Salarpour